The weight of fruit in a plant is determined additively by 3 genes. A plant having all recessive alleles has a fruit weight of 2 g. A plant having all dominant alleles has a fruit weight of 8 g. Determine the range of weights (minimum and maximum) for the offspring of the following cross: AA Bb CC x aa Bb Cc

  1. 4 – 7 grams
  2. 2 – 8 grams
  3. 4 – 8 grams
  4. 2 – 5 grams
Answer #1

1. Answer: Option A is correct
Explanation:
It is given that fruit weight is determined by three different genes, each gene with two additive alleles.
aabbcc = 2 g
AABBCC = 8 g
Contribution from each dominant allele = (8-2)/6
= 1 g

Parental cross: AABb CC X aaBbCc
Maximum fruit length = AaBBCC = 2+5 = 7 g
Minimum fruit length = AabbCc = 2+2 = 4 g

A positive rating is highly appreciated. Thank you